OpenStudy (anonymous):

Question 1 Solve: –5x = 30 6 –6 25 35 Question 2 Simplify: (9 – 8)2 × (24 – 6 3) 6 12 22 44 Question 3 Solve: 2x – 9 = 15 3 4 12 48 Question 4 Solve: –2x – 15 = 6x + 9 3 –3 –1 6 Question 5 Solve: –4(2x – 1) = –6(x + 2) – 2 –5 5 9 –9

OpenStudy (anonymous):

1) x is 6 because you need to divide 30 by 5 so you can get x by itself

OpenStudy (anonymous):

2) (9 - 8)²(24 - 63) = 1²(- 39) = - 39...The answer is -39

OpenStudy (anonymous):

3) okay, subtract 9 from both sides -2x = 15-9 -2x = 6 x = -3

OpenStudy (anonymous):

1;x=-6 3;x=12 4; x=-4 5;9

OpenStudy (anonymous):

4) -2x-15=6x+9 -8x=24 -x=3 x=-3

OpenStudy (anonymous):

-4(2x-1)= -6(x+2)-2 -4(2x)-4(-1)= -6(x)-6(2)-2 -8x+4=-6x-12-2 -8x+4=-6x-14 -8x=-6x-18 8x=6x+18 2x=18 x=9

OpenStudy (anonymous):

2;= (9 - 8)²(24 - 63) = 1²(- 39) = - 39 Answer: - 39

OpenStudy (anonymous):

1;